# Adaptive Collaborative Correlation Learning-based Semi-Supervised Multi-Label Feature Selection

### v1

@@ 2 Related Work
Unlike the methods in the first category, the second category explores label correlations and constructs the feature selection model directly on multi-label data. MIFS selects discriminative features by factorizing the label matrix to reveal label correlations and diminish the impact of imperfect label information [13]. In GLFS [17], the LASSO-based sparse learning is used to select label-specific features that preserve groups, enabling the simultaneous exploration of label-group and instance-group correlations. However, these methods assume that all samples are labeled, which may not hold true in real-world scenarios. More often, only a subset of samples are labeled. Multi-label feature selection within the semi-supervised learning paradigm is proposed to address this issue. In LEDA [22], Guo et al. devised a semi-supervised multi-label propagation method to capture information from both labeled and unlabeled data simultaneously. Subsequently, the acquired label information is incorporated into an extended linear discriminant analysis module to select informative features. Xu et al. developed a semi-supervised multi-label feature selection algorithm that utilizes the probabilistic neighborhood similarities to learn the correlations between features and maintains consistent correlation information within the label space [27]. To explore the local manifold structures of both feature and label spaces in semi-supervised multi-label data, SMLFS [28] uses a local logistic regression model that incorporates feature graph regularization and label graph regularization. In addition, ${\ell_{2,p}}$-norm is applied to the regression coefficient matrix to select important feature dimensions. In [29], Li et al. proposed a semi-supervised multi-label feature selection method that simultaneously utilizes label and sample correlations. By imposing binary hash constraints on the spectral embedding model, SFS-BLL [30] identifies pseudo-labels for unlabeled data and utilizes a self-weighted sparse regression model to select discriminative features, leveraging both existing labels and learned pseudo-labels.
@@ 3.2 Formulation of Access-MFS
To select the discriminative and uncorrelated features, we propose an extended uncorrelated constraint on $\mathbf{W}$, denoted as $\mathbf{W}^{T}\mathbf{R}\mathbf{W}=\mathbf{I}$. Here, $\mathbf{R}$ is defined as $\mathbf{X}\mathbf{H}\mathbf{X}^{T}+\lambda\mathbf{D}+\theta\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}$, where the matrix $\mathbf{D}\in\mathbb{R}^{d\times d}$ is diagonal, with the $i$th diagonal entry given by $\mathbf{D}(i,i)=\frac{1}{2\sqrt{\|\mathbf{w}_{i.}\|_{2}^{2}+\epsilon}}$ ($\epsilon$ is a small constant to avoid division by 0). Additional, $\mathbf{L}_{s}=\mathbf{D}_{s}-\mathbf{S}$ represents the Laplacian matrix of instance similarity matrix $\mathbf{S}$, $\mathbf{D}_{s}$ is the degree matrix of $\mathbf{S}$, and $\theta$ is a trade-off parameter. The uncorrelated constraint we proposed consists of three terms obtained by expanding $\mathbf{R}$. The first term aims to promote the orthonormality of the projected dimensions, which makes the low-dimensional instances to be uncorrelated. The second term is added to the first term to avoid the singularity of $\mathbf{X}\mathbf{H}\mathbf{X}^{T}$. The third term is advantageous in reducing the variance among samples within the same neighborhood under the graph structure. When $\lambda=\theta=0$, the proposed constraint reduces to the traditional uncorrelated constraint. The proposed constraint is beneficial to select discriminative features that are also uncorrelated. Furthermore, it can simplify the optimization process of $\mathbf{W}$ and transforms $\mathbf{W}$ into a closed-form solution, as demonstrated in the optimization procedure.
@@ 4.1 Update $\mathbf{W}$ by Fixing Other Variables
Input: Data matrix $\mathbf{X}$, centering matrix $\mathbf{H}$, predicted label matrix $\mathbf{F}$ and Laplacian matrix $\mathbf{L}_{s}$, parameters: $\lambda$ and $\theta$.
@@ 14            Update $\mathbf{W}=\mathbf{R}^{-\frac{1}{2}}\mathbf{A}$;
4.2 Update $\mathbf{F}$ by Fixing Other Variables
$\mathbf{Q}\mathbf{F}+\mu\mathbf{F}\mathbf{L}_{p}=\mathbf{C}$ (17)
@@ 4.4 Update $\mathbf{P}$ by Fixing Other Variables
5 Convergence and Complexity Analysis
5.1 Convergence Analysis
Theorem 1.
Proof.
Let $\varUpsilon(\mathbf{W}^{(t)},\mathbf{D}^{(t)})=\|\mathbf{H}\mathbf{X}^{T}\mathbf{W}^{(t)}-\mathbf{H}\mathbf{F}\|_{F}^{2}+\lambda\operatorname{Tr}({\mathbf{W}^{(t)}}^{T}\mathbf{D}^{(t)}\mathbf{W}^{(t)})+\theta\operatorname{Tr}({\mathbf{W}^{(t)}}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W}^{(t)})$ represent the objective value in the $t$th iteration of Eq. (10). According to [35], during the alternate updating of $\mathbf{W}$ and $\mathbf{D}$ using Algorithm 1, the following inequalities must hold for the ($t$+1)th iteration:
$\varUpsilon(\mathbf{W}^{(t+1)},\mathbf{D}^{(t)})\leq\varUpsilon(\mathbf{W}^{(t)},\mathbf{D}^{(t)})$ (26)
Besides, by means of the property $\|\mathbf{u}\|_{2}-\frac{\|\mathbf{u}\|_{2}^{2}}{2\|\mathbf{v}\|_{2}}\leq\|\mathbf{v}\|_{2}-\frac{\|\mathbf{v}\|_{2}^{2}}{2\|\mathbf{v}\|_{2}}$ for any two nonzero vectors $\mathbf{u}$ and $\mathbf{v}$ as proposed in [34], we have
$\|\mathbf{w}^{(t+1)}_{i.}\|_{2}-\frac{\|\mathbf{w}^{(t+1)}_{i.}\|_{2}^{2}}{2\|\mathbf{w}^{(t)}_{i.}\|_{2}}\leq\|\mathbf{w}^{(t)}_{i.}\|_{2}-\frac{\|\mathbf{w}^{(t)}_{i.}\|_{2}^{2}}{2\|\mathbf{w}^{(t)}_{i.}\|_{2}}$
$\Rightarrow$ $\lambda\sum_{i=1}^{d}\|\mathbf{w}^{(t+1)}_{i.}\|_{2}-\lambda\sum_{i=1}^{d}\frac{\|\mathbf{w}^{(t+1)}_{i.}\|_{2}^{2}}{2\|\mathbf{w}^{(t)}_{i.}\|_{2}}\leq$
$\lambda\sum_{i=1}^{d}\|\mathbf{w}^{(t)}_{i.}\|_{2}-\lambda\sum_{i=1}^{d}\frac{\|\mathbf{w}^{(t)}_{i.}\|_{2}^{2}}{2\|\mathbf{w}^{(t)}_{i.}\|_{2}}$
$\Rightarrow$ $\lambda\|\mathbf{W}^{(t+1)}\|_{2,1}-\lambda\operatorname{Tr}({\mathbf{W}^{(t+1)}}^{T}\mathbf{D}^{(t)}\mathbf{W}^{(t+1)})\leq$
$\lambda\|\mathbf{W}^{(t)}\|_{2,1}-\lambda\operatorname{Tr}({\mathbf{W}^{(t)}}^{T}\mathbf{D}^{(t)}\mathbf{W}^{(t)}).$ (27)
By combing the inequalities (26) and (27), it is easy to obtain
$\|\mathbf{H}\mathbf{X}^{T}\mathbf{W}^{(t+1)}-\mathbf{H}\mathbf{F}\|_{F}^{2}+\lambda\|\mathbf{W}^{(t+1)}\|_{2,1}$ (28)
$+\theta\operatorname{Tr}({\mathbf{W}^{(t+1)}}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W}^{(t+1)})$
$\leq\|\mathbf{H}\mathbf{X}^{T}\mathbf{W}^{(t)}-\mathbf{H}\mathbf{F}\|_{F}^{2}+\lambda\|\mathbf{W}^{(t)}\|_{2,1}$
$+\theta\operatorname{Tr}({\mathbf{W}^{(t)}}^{T}\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}\mathbf{W}^{(t)}).$
Hence, the objective function in Eq. (9) monotonically decreases by Algorithm 2 in each iteration. ∎
As Eq. (15) is equivalent to Eq. (16), we only need to demonstrate that the objective function in Eq.(16) with respect to ${\mathbf{F}}$ is a monotonically decreasing function. It is easy to verify that the matrix $\mathbf{Q}$ and the Laplacian matrix $\mathbf{L}_{p}$ are both positive semidefinite. Consequently, the quadratic forms $\mathbf{F}^{T}\mathbf{Q}\mathbf{F}-2\mathbf{F}^{T}\mathbf{C}$ and $\mathbf{F}\mathbf{L}_{p}\mathbf{F}^{T}$ are convex. Hence, the objective function in Eq.(16) is convex. According to [36], we can easily obtain that the objective function in Eq.(16) monotonically decrease by using the updating rules in Algorithm 2. Additionally, the objective functions in Eqs. (18) and (22) are convex w.r.t. $\mathbf{S}$ and $\mathbf{P}$, and possess closed-form solutions, ensuring the convergence of the updates for $\mathbf{S}$ and $\mathbf{P}$, respectively. Therefore, we can conclude that the objective function in Eq.(8) monotonically decreases with each iteration of Algorithm 2. Moreover, the convergence behavior of Algorithm 2 is further demonstrated through experiments.

### v3

@@ 2 Related Work
Unlike the methods in the first category, the second category explores label correlations and constructs the feature selection model directly on multi-label data. MIFS selects discriminative features by factorizing the label matrix to reveal label correlations and diminish the impact of imperfect label information [13]. In GLFS [17], the LASSO-based sparse learning is used to select label-specific features that preserve groups, enabling the simultaneous exploration of label-group and instance-group correlations. However, these methods assume that all samples are labeled, which may not hold true in real-world scenarios. More often, only a subset of samples are labeled. Multi-label feature selection within the semi-supervised learning paradigm is proposed to address this issue. In LEDA [22], Guo et al. devised a semi-supervised multi-label propagation method to capture information from both labeled and unlabeled data simultaneously. Subsequently, the acquired label information is incorporated into an extended linear discriminant analysis module to select informative features. To explore the local manifold structures of both feature and label spaces in semi-supervised multi-label data, SMLFS [27] uses a local logistic regression model that incorporates feature graph regularization and label graph regularization. In addition, ${\ell_{2,p}}$-norm is applied to the regression coefficient matrix to select important feature dimensions. In [28], Li et al. proposed a semi-supervised multi-label feature selection method that simultaneously utilizes label and sample correlations. By imposing binary hash constraints on the spectral embedding model, SFS-BLL [29] identifies pseudo-labels for unlabeled data and utilizes a self-weighted sparse regression model to select discriminative features, leveraging both existing labels and learned pseudo-labels.
@@ 3.2 Formulation of Access-MFS
To select the discriminative and uncorrelated features, we propose an extended uncorrelated constraint on $\mathbf{W}$, denoted as $\mathbf{W}^{T}\mathbf{R}\mathbf{W}=\mathbf{I}$. Here, $\mathbf{R}$ is defined as $\mathbf{X}\mathbf{H}\mathbf{X}^{T}+\lambda\mathbf{D}+\theta\mathbf{X}\mathbf{L}_{s}\mathbf{X}^{T}$, where the matrix $\mathbf{D}\in\mathbb{R}^{d\times d}$ is diagonal, with the $i$th diagonal entry given by $\mathbf{D}(i,i)=\frac{1}{2\sqrt{\|\mathbf{w}_{i.}\|_{2}^{2}+\epsilon}}$ ($\epsilon$ is a small constant to avoid division by 0). Additional, $\mathbf{L}_{s}=\mathbf{D}_{s}-\mathbf{S}$ represents the Laplacian matrix of instance similarity matrix $\mathbf{S}$, $\mathbf{D}_{s}$ is the degree matrix of $\mathbf{S}$, and $\theta$ is a trade-off parameter. The uncorrelated constraint we proposed consists of three terms obtained by expanding $\mathbf{R}$. The first term aims to promote the orthonormality of the projected dimensions, which makes the low-dimensional instances to be uncorrelated. The second term is added to the first term to avoid the singularity of $\mathbf{X}\mathbf{H}\mathbf{X}^{T}$. The third term is advantageous in reducing the variance among samples within the same neighborhood under the graph structure. When $\lambda=\theta=0$, the proposed constraint reduces to the traditional uncorrelated constraint. The proposed constraint is beneficial to select discriminative features that are also uncorrelated.
@@ 4.1 Update $\mathbf{W}$ by Fixing Other Variables
4.2 Update $\mathbf{F}$ by Fixing Other Variables
$\mathbf{Q}\mathbf{F}+\mu\mathbf{F}\mathbf{L}_{p}=\mathbf{C}$ (17)
Input: Data matrix $\mathbf{X}$, centering matrix $\mathbf{H}$, predicted label matrix $\mathbf{F}$, Laplacian matrix $\mathbf{L}_{s}$, and the parameters $\lambda$ and $\theta$.
@@ 4             Update $\mathbf{W}$ by using Algorithm 1;
5 Convergence and Complexity Analysis
5.1 Convergence Analysis
Theorem 1.
We can prove this theorem based on the property in [33] and the result in [34]. Due to space constraints, the detailed proof of Theorem 1 is provided in the supplementary material.
Besides, as Eq. (15) is equivalent to Eq. (16), we only need to demonstrate that the objective function in Eq.(16) with respect to ${\mathbf{F}}$ is a monotonically decreasing function. It is easy to verify that the matrix $\mathbf{Q}$ and the Laplacian matrix $\mathbf{L}_{p}$ are both positive semidefinite. Consequently, the quadratic forms $\mathbf{F}^{T}\mathbf{Q}\mathbf{F}-2\mathbf{F}^{T}\mathbf{C}$ and $\mathbf{F}\mathbf{L}_{p}\mathbf{F}^{T}$ are convex. Hence, the objective function in Eq.(16) is convex. According to [35], we can easily obtain that the objective function in Eq.(16) monotonically decrease by using the updating rules in Algorithm 2. Additionally, the objective functions in Eqs. (18) and (22) are convex w.r.t. $\mathbf{S}$ and $\mathbf{P}$, and possess closed-form solutions, ensuring the convergence of the updates for $\mathbf{S}$ and $\mathbf{P}$, respectively. Therefore, we can conclude that the objective function in Eq.(8) monotonically decreases with each iteration of Algorithm 2. Moreover, the convergence behavior of Algorithm 2 is further demonstrated through experiments.